For the hydrolysis of a salt of a weak acid and a weak base, show that
a b
w
K K
K
h
K 
(5)
1
Expert's answer
2020-02-17T07:10:46-0500

Solution.

For weak acid:

"Ka = \frac{[H^+][A^-]}{[HA]}"

For weak base:

"Kb = \frac{[B^+][OH^-]}{[BOH]}"

Final reaction:

"A^- + B^+ + H2O = HA + BOH"

"Kh = \frac{[HA][BOH]}{[A^-][B^+]}"

So, we get:

"Kh = \frac{Kw}{Ka \times Kb}"

Because:

"Kw = [H^+][OH^-]"

Answer:

"Kh = \frac{Kw}{Ka \times Kb}"